OpenStudy (anonymous):

A cooler contains 6 cans of soda; 2 colas, 3 orange, and 1 cherry. Two cans are selected at random without replacement. Find the probability that at least one can is cola.

OpenStudy (anonymous):

first you have to add them all up, so go ahead and do that

OpenStudy (anonymous):

you should get 6. so they want the probability that one of the two cans pulled out is cola. so 2 out of the 6 sodas are cola. so the probability is 2 out of 6.

OpenStudy (anonymous):

2/6= 1/3?

OpenStudy (anonymous):

yes you can simplify

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

no problem:)

OpenStudy (raden):

the probably events are (1 cola and 1 orange) + (1 cola and 1 cherry) + (2 colas's) so, the probability of that events = P(1Co and 1O) + P(1Co and 1Ch) + P(2cola's) = 2/6 * 3/5 + 2/6 * 1/5 + 2/6 * 1/5 = 6/30 + 2/30 + 2/30 = 10/30 = 1/3